OpenStudy (johanna1736):

Solve: negative 1 over 2 x + 3 = –x + 7

OpenStudy (johanna1736):

Please help

OpenStudy (fortytherapper):

\[-\frac{ 1 }{ 2 }x+3=-x+7\] Is that it?

OpenStudy (johanna1736):

8 –8 2

OpenStudy (johanna1736):

yes.

OpenStudy (johanna1736):

Answer choices: A. 1/2 B. 8 C. 2 D. -8

OpenStudy (fortytherapper):

Let's turn that -x into a fraction to make it look easier \[-\frac{ 1 }{ 2 }x+3=-\frac{ 2 }{ 2 }x+7\] So, how could we get the 3 over to the other side of the = sign?

OpenStudy (anonymous):

take terms of x one one side and numbers on other \[x-\frac{ x }{ 2 }=7-3\]

OpenStudy (johanna1736):

We would subtract the three?

OpenStudy (fortytherapper):

Yeah, ash has already done those steps, so you can combine the like terms in his answer

OpenStudy (johanna1736):

how would we do that

OpenStudy (fortytherapper):

Do them separately to make it easier \[x-\frac{ x }{ 2 }=?\] \[7-3=?\]

OpenStudy (johanna1736):

x/2 an 4

OpenStudy (fortytherapper):

Now we can replace that with \[\frac{ x }{ 2 } = 4\]

OpenStudy (johanna1736):

Then we multiply both sides by 2?

OpenStudy (fortytherapper):

Correct

OpenStudy (johanna1736):

soo x=4

OpenStudy (johanna1736):

8.

OpenStudy (johanna1736):

x=8?

OpenStudy (fortytherapper):

Yeah, 8, good job

OpenStudy (johanna1736):

thx

OpenStudy (anonymous):

@FortyTheRapper you are a good teacher man.